# Efficient quantum Gibbs samplers with Kubo–Martin–Schwinger detailed balance condition

### v2

@@ 3.2. Choice of the weighting function $q(\nu)$
    Here $u(\nu)e^{-\frac{\nu}{4}}=e^{-\frac{(\nu+1)^{2}-1}{8}}$ is a Gevrey function of order $s_{u}=\frac{1}{2}$ by [HR19, Proposition B.1] and the $L^{1}$-integrability of its derivative is straightforward. So we still have $s=\max\{s_{u},s_{w}\}=s_{w}$. Setting $S=\mathcal{O}(1/\beta)$, there holds
@@
In this section, we discuss the efficient simulation of the Lindblad equation in (3.9). For simplicity, we assume that $q^{a}=q$, $f^{a}=f$, and $g^{a}=g$ and $q$ satisfies 15. Our construction uses the block encoding input model and the linear combination of unitaries (see Appendices A and B). We also assume $\max_{a\in\mathcal{A}}\|A^{a}\|\leq 1$ to ensure that efficient block encodings of $A^{a}$ are available (see Eq. 3.23).
@@
According to 16, in the following discussion, we set the integer $\mathfrak{m}$ large enough and consider the quadrature points $\{t_{m}\}^{2M-1}_{m=0}$ ($M=2^{\mathfrak{m}-1}$) as in Eq. 3.13 such that Eq. 3.14 holds. Without loss of generality, we assume $|\mathcal{A}|=2^{\mathfrak{a}}$ with $\mathfrak{a}\in\mathbb{N}_{+}$.
@@
    (3.25) $\textbf{Prep}_{\mathcal{A}}=\mathrm{H}^{\otimes\mathfrak{a}}:\ket{0^{\mathfrak{a}}}=\frac{1}{\sqrt{|\mathcal{A}|}}\sum_{a\in\mathcal{A}}\ket{a}\,,$
@@
    Here the block encoding factor $Z_{f}:=\sum^{2M-1}_{m=0}|f(t_{m})|\tau$ is bounded by 32 with Eq. D.3:
    (3.28) $Z_{f}=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log(\beta A_{u}+A_{w})\right)\,.$
@@
    (3.29) $\textbf{Prep}_{g}:\ket{0^{\mathfrak{m}}}\ket{0^{\mathfrak{m}}}=\frac{1}{\sqrt{Z_{g}}}\sum^{2M-1}_{n,m=0}\sqrt{g(t_{n},t_{m})\tau^{2}}\ket{t_{n}}\ket{t_{m}}\,,$
@@
    (3.30) $\textbf{Prep}_{\overline{g}}:\ket{0^{\mathfrak{m}}}\ket{0^{\mathfrak{m}}}=\frac{1}{\sqrt{Z_{g}}}\sum^{2M-1}_{n,m=0}\sqrt{\overline{g(t_{n},t_{m})}\tau^{2}}\ket{t_{n}}\ket{t_{m}}\,,$
    with the block encoding factor $Z_{g}=\sum^{2M-1}_{n,m=0}|g(t_{n},t_{m})|\tau^{2}$ bounded as follows, by 32 with Eq. D.4,
    (3.31) $Z_{g}=\mathcal{O}\left((A_{q}^{2}C_{1,u}^{2}+A_{q}^{4}S^{2}A_{w}^{2}+A_{q}^{3}C_{1,u}SA_{w}+A_{q}^{4}\beta S)\log^{2}(\beta A_{u}+A_{w})\right)\,.$
Using these oracles, according to 16, we can apply LCU with (3.23), (3.24), (3.26), and (3.27) to construct a
$\left(Z_{f}Z_{\mathcal{A}},\mathfrak{m}+\mathfrak{b},C_{f}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta A_{u}+A_{w})^{1/s}e}\right)\right)$
@@
Here $C_{f}$ is defined in 16. The circuit of $U_{L}$ can be found in Fig. 2. In (3.32), the total Hamiltonian simulation time required by one query to $U_{L}$ is $\mathcal{O}\left(M\tau\right)$.
@@
$\left(Z_{g}Z^{2}_{\mathcal{A}}\sqrt{|\mathcal{A}|},2\mathfrak{m}+\mathfrak{a}+\mathfrak{b},C_{g}|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\right)\right)$
@@
(3.33) $U_{G}=$ $\underbrace{\left(\mathrm{I}^{\otimes 2}_{\mathfrak{m}}\otimes\textbf{Prep}_{\mathcal{A}}\otimes\mathrm{I}_{\mathfrak{b}}\otimes\mathrm{I}_{n}\right)^{\dagger}\cdot\left(\textbf{Prep}^{\dagger}_{\overline{g}}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{b}}\otimes\mathrm{I}_{n}\right)\cdot}_{\text{Two layers of LCU prepare oracles}}$
$\underbrace{\mathrm{I}_{\mathfrak{m}}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{b}}\otimes U_{H}^{\dagger}}_{\exp(iHt_{n})}\cdot\underbrace{(\mathrm{I}_{\mathfrak{m}}^{\otimes 2}\otimes U_{\mathcal{A}})}_{A^{a}}\cdot\underbrace{\mathrm{I}_{\mathfrak{m}}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{b}}\otimes U_{H}}_{\exp(-iHt_{n})}\cdot\underbrace{\mathrm{I}_{\mathfrak{m}}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{b}}\otimes U^{\dagger}_{H}}_{\exp(iHt_{m})}\cdot\underbrace{(\mathrm{I}_{\mathfrak{m}}^{\otimes 2}\otimes U_{\mathcal{A}})}_{A^{a}}\cdot\underbrace{\mathrm{I}_{\mathfrak{m}}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{b}}\otimes U_{H}}_{\exp(-iHt_{m})}$
$\underbrace{\cdot\left(\textbf{Prep}_{g}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{b}}\otimes\mathrm{I}_{n}\right)\cdot\left(\mathrm{I}^{\otimes 2}_{\mathfrak{m}}\otimes\textbf{Prep}_{\mathcal{A}}\otimes\mathrm{I}_{\mathfrak{b}}\otimes\mathrm{I}_{n}\right)}_{\text{Two layers of LCU prepare oracles}}\,.$
Here $C_{g}$ is defined in 16. We note that $U_{H}$ for $\exp(-iHt_{m})$ and $\exp(-iHt_{n})$ are acting on different time registers. The circuit of $U_{G}$ is given in Fig. 2. In (3.33), one query to $U_{G}$ still requires $\mathcal{O}\left(M\tau\right)$ total Hamiltonian simulation time.
@@
Figure 2. Quantum circuits for block encodings of $\{L_{a}\}_{a\in\mathcal{A}}$ (top) and $G$ (bottom). The simulation of $e^{iHt_{n}}$ and $e^{-iHt_{m}}$ involved in the block encoding $U_{G}$ of $G$ can be combined into a single step $e^{iH(t_{n}-t_{m})}$ via Hamiltonian simulation controlled simultaneously by the top two registers.
@@
Now, we are ready to give the simulation cost of our method as follows.
@@
Assume access to weighting functions $\{q^{a}\}$ satisfying 15 with any $s>1$, block encodings $U_{\mathcal{A}}$ in Eq. 3.23, controlled Hamiltonian simulation $U_{H}$ in Eq. 3.24, and prepare oracles for filtering functions $\{f^{a}\}$ and $\{g^{a}\}$ in (3.26)–(3.30). The Lindbladian evolution (3.9) can be simulated up to time $t_{\rm mix}$ with an $\epsilon$-diamond distance, and the total Hamiltonian simulation time is
$\widetilde{\mathcal{O}}\left(C_{q}t_{\rm mix}\beta^{2}S|\mathcal{A}|^{2}\log^{1+s}\left(1/\epsilon\right)\right)\,,$
@@
$C_{q}:=A_{q}^{2}C_{1,u}^{2}/\beta S+A_{q}^{4}A_{w}^{2}S/\beta+A_{q}^{3}C_{1,u}A_{w}/\beta+A_{q}^{4}\,.$
@@
$\widetilde{\mathcal{O}}\left(\log(A_{q}S)+\log^{2}(t_{\rm mix}|\mathcal{A}|/\epsilon)+\log^{2}\left(C_{q}\beta S\right)+\log A_{u}\right)\,.$
number of additional ancilla qubits for the prepare oracles and simulation. The $\widetilde{\mathcal{O}}$ absorbs a constant only depending on $s$ and subdominant polylogarithmic dependencies on parameters $t_{\rm mix}$, $|\mathcal{A}|$, $S$, $A_{q}$, $A_{u}$, $A_{w}$, and $\beta$.
Remark 20.
In the above theorem, the total Hamiltonian simulation time scales quadratically in $\beta$, a complexity seemingly less favorable compared to the one in [CKG23], which is only linearly dependent in $\beta$ (see Table 1). We will provide a detailed explanation of this distinction in Section 4. However, at first glance, we note that if $q$ is chosen as a concentrated filtering function with support width $S=\mathcal{O}(1/\beta)$, one can readily recover a linear scaling of the total Hamiltonian simulation time in $\beta$. Although this choice offers a better theoretical simulation complexity, it confines energy transitions to a narrow range, potentially prolonging the mixing time. Balancing this trade-off by selecting $S$ appropriately constitutes an intriguing avenue for future research.
Proof.
For simplicity, we let
(3.35) ${\rm C}_{q,\beta}:=A_{q}^{2}C_{1,u}^{2}+A_{q}^{4}S^{2}A_{w}^{2}+A_{q}^{3}C_{1,u}SA_{w}+A_{q}^{4}\beta S\,.$
By 18 with estimates in Eqs. 3.28 and 3.31, we have
(3.36) $\|\mathcal{L}\|_{\rm be}$ $=Z_{g}Z^{2}_{\mathcal{A}}\sqrt{|\mathcal{A}|}+\frac{1}{2}Z^{2}_{f}Z^{2}_{\mathcal{A}}|\mathcal{A}|$
$=\mathcal{O}\left({\rm C}_{q,\beta}\log^{2}(\beta A_{u}+A_{w})|\mathcal{A}|\right)\,,$
Recalling 16 and 32, we set a truncation time
(3.37) $T=\Theta\left((\beta A_{u}+A_{w})\log^{s}\left(\frac{t_{\rm mix}\|\mathcal{L}\|_{\rm be}(C_{f}+C_{g}|\mathcal{A}|)}{\epsilon}\right)\right)\,,$
and the step size
(3.38) $\tau=\Theta\left(1/A_{q}^{4}T^{2}S^{3}\right)\,,$
$(C_{f}+C_{g}|\mathcal{A}|)\exp\left(-\frac{sT^{1/s}}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\right)\leq\epsilon/{t_{\rm mix}}\|\mathcal{L}\|_{\rm be}\,.$
Note that one query to the block encoding of $L_{a}$ requires one query to $U_{L}$, by 18 and Eq. 3.36, the simulation requires
$=$ $\,\widetilde{\mathcal{O}}\left(t_{\rm mix}{\rm C}_{q,\beta}\log^{2}(\beta A_{u}+A_{w})|\mathcal{A}|^{2}\log\left(1/\epsilon\right)\right)$
queries to $U_{L}$. Similarly, we need
$\widetilde{\mathcal{O}}\left(t_{\rm mix}{\rm C}_{q,\beta}\log^{2}(\beta A_{u}+A_{w})|\mathcal{A}|\log\left(1/\epsilon\right)\right)$
$\widetilde{\mathcal{O}}\left(t_{\rm mix}{\rm C}_{q,\beta}(\beta A_{u}+A_{w})|\mathcal{A}|^{2}\log^{1+s}\left(1/\epsilon\right)\right)=\widetilde{\mathcal{O}}\left(C_{q}t_{\rm mix}\beta^{2}S|\mathcal{A}|^{2}\log^{1+s}\left(1/\epsilon\right)\right)\,.$
Next, we consider the number of extra ancilla qubits, again by 18. We first note
$\log(t_{\rm mix}\left\lVert\mathcal{L}_{\rm be}\right\rVert/\epsilon)=\widetilde{\mathcal{O}}\left(\log(t_{\rm mix}|\mathcal{A}|/\epsilon)+\log{\rm C}_{q,\beta}+\log\log A_{u}\right)$
and hence there holds
$=$ $\widetilde{\mathcal{O}}\left(\log^{2}(t_{\rm mix}|\mathcal{A}|/\epsilon)+(\log{\rm C}_{q,\beta}+\log\log A_{u})^{2}\right)\,.$
For the preparation oracles, by Eqs. 3.38 and 3.37, we find
$M=\Theta(A_{q}^{4}T^{3}S^{3})=\Theta\left(A_{q}^{4}S^{3}(\beta A_{u}+A_{w})^{3}\log^{3s}\left(\frac{t_{\rm mix}\|\mathcal{L}\|_{\rm be}(C_{f}+C_{g}|\mathcal{A}|)}{\epsilon}\right)\right)$
This, along with Eqs. 3.16 and 3.18, implies that we need
$\mathfrak{m}=\mathcal{O}\left(\log(A_{q}S)+\log(\beta A_{u}+A_{w})+\log\left(t_{\rm mix}\|\mathcal{L}\|_{\rm be}|\mathcal{A}|/\epsilon\right)\right)$

### v4

@@ 3.2. Choice of weighting function $q(\nu)$ and $\kappa(\nu)$
    Here $u(\nu)e^{-\frac{\nu}{4}}=e^{-\frac{(\nu+1)^{2}-1}{8}}$ is a Gevrey function of order $1$ by [HR19, Proposition B.1] and the $L^{1}$-integrability of its derivative is straightforward. Setting $S=\mathcal{O}(1/\beta)$, there holds
@@
In this section, we discuss the efficient simulation of the Lindblad equation in (3.13). For simplicity, we assume that $q^{a}=q$ satisfies 13 and then $f^{a}=f$. Our construction uses the block encoding input model and the linear combination of unitaries (see Appendices A and B). We also assume $\max_{a\in\mathcal{A}}\|A^{a}\|\leq 1$ to ensure that efficient block encodings of $A^{a}$ are available (see Eq. 3.24).
@@
According to 15, in the following discussion, we set the integer $\mathfrak{m}$ large enough and consider the quadrature points $\{t_{m}\}^{2M-1}_{m=0}$ ($M=2^{\mathfrak{m}-1}$) as in Eq. 3.18 such that Eq. 3.19 holds. Without loss of generality, we assume $|\mathcal{A}|=2^{\mathfrak{a}}$ with $\mathfrak{a}\in\mathbb{N}_{+}$.
@@
    (3.26) $\textbf{Prep}_{\mathcal{A}}=\mathrm{H}^{\otimes\mathfrak{a}}\ket{0^{\mathfrak{a}}}=\frac{1}{\sqrt{|\mathcal{A}|}}\sum_{a\in\mathcal{A}}\ket{a}\,,$
@@
    Here the block encoding factor $Z_{f}:=\sum^{2M-1}_{m=0}|f(t_{m})|\tau$ is bounded by 33 with Eq. D.3:
    (3.29) $Z_{f}=\mathcal{O}\left((\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})\log((\beta\xi_{u}+\xi_{w}/S)\max\{S,1\})\right)\,.$
@@
    (3.30) $\textbf{Prep}_{g}:\ket{0^{\mathfrak{m}}}=\frac{1}{\sqrt{Z_{g}}}\sum^{2M-1}_{m=0}\sqrt{g(t_{m})\tau}\ket{t_{m}}\,,$
@@
    (3.31) $\textbf{Prep}_{\overline{g}}:\ket{0^{\mathfrak{m}}}=\frac{1}{\sqrt{Z_{g}}}\sum^{2M-1}_{m=0}\sqrt{\overline{g(t_{m})}\tau}\ket{t_{m}}\,,$
    with the block encoding factor $Z_{g}=\sum^{2M-1}_{m=0}|g(t_{m})|\tau$ bounded by 33 with Eq. D.4:
    (3.32) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Z_{g}=\mathcal{O}\left(\xi_{q}(1+\log((\beta+\xi_{w}/S)\max\{S,1\}))\right)\,.}$
Using these oracles, according to 15, we can apply LCU with (3.24), (3.25), (3.27), and (3.28) to construct a
$\left(Z_{f}Z_{\mathcal{A}},\mathfrak{m}+\mathfrak{a}+\mathfrak{b},C_{f}S|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\right)\right)$
@@
Here $C_{f}$ is defined in 15. The circuit of $U_{L}$ can be found in Fig. 2. In (3.33), the total Hamiltonian simulation time required by one query to $U_{L}$ is $\mathcal{O}\left(M\tau\right)$.
@@
$\left(Z_{g}Z^{2}_{\mathcal{A}}|\mathcal{A}|,3\mathfrak{m}+\mathfrak{a}+2\mathfrak{b},C_{g}S|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2e(\beta+\xi_{w}/4S)^{1/s}}\right)\right)$
@@
(3.34) $U_{G}=$ $\underbrace{\left(\mathrm{I}_{\mathfrak{m}}\otimes\textbf{Prep}_{\mathcal{A}}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\otimes\mathrm{I}_{n}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\right)^{\dagger}\cdot\left(\textbf{Prep}^{\dagger}_{\overline{g}}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{m}+\mathfrak{b}}\otimes\mathrm{I}_{n}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\right)\cdot}_{\text{Two layers of LCU prepare oracles}}$
$\underbrace{\mathrm{I}_{\mathfrak{a}+2\mathfrak{m}+\mathfrak{b}}\otimes U_{H}^{\dagger}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}}_{\exp(iHt_{n})}\cdot\underbrace{\left(\mathrm{I}_{\mathfrak{m}}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\otimes U^{\dagger}_{L}\right)\left(\mathrm{I}_{\mathfrak{m}}\otimes U_{L}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\right)}_{\mathinner{|{a}\rangle\!\langle{a}|}\otimes L^{\dagger}_{a}L_{a}}\cdot\underbrace{\mathrm{I}_{\mathfrak{a}+2\mathfrak{m}+\mathfrak{b}}\otimes U_{H}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}}_{\exp(-iHt_{n})}$
$\underbrace{\cdot\left(\textbf{Prep}_{g}\otimes\mathrm{I}_{\mathfrak{a}+\mathfrak{m}+\mathfrak{b}}\otimes\mathrm{I}_{n}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\right)\cdot\left(\mathrm{I}_{\mathfrak{m}}\otimes\textbf{Prep}_{\mathcal{A}}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\otimes\mathrm{I}_{n}\otimes\mathrm{I}_{\mathfrak{m}+\mathfrak{b}}\right)}_{\text{Two layers of LCU prepare oracles}}\,.$
Here $C_{g}$ is defined in 15. We note that $U_{H}$ for $\exp(-iHt_{m})$ and $\exp(-iHt_{n})$ are acting on different time registers. The circuit of $U_{G}$ is given in Fig. 2. In (3.34), one query to $U_{G}$ still requires $\mathcal{O}\left(M\tau\right)$ total Hamiltonian simulation time.
@@
Figure 2. Quantum circuits for block encodings of $\{L_{a}\}_{a\in\mathcal{A}}$ (top) and $G$ (bottom).
@@
Now, we are ready to give the simulation cost of our method as follows, where we have assumed $S\geq 1$ for simplicity as in 15. The proof is postponed to Appendix D.
@@
Assume access to weighting functions $\{q^{a}\}$ satisfying 13 with $s>1$, block encodings $U_{\mathcal{A}}$ in Eq. 3.24, controlled Hamiltonian simulation $U_{H}$ in Eq. 3.25, and prepare oracles for filtering functions $\{f^{a}\}$ and $g$ in (3.27)–(3.31). The Lindbladian evolution (3.13) with $\beta>0$ can be simulated up to time $t_{\rm mix}$ within $\epsilon$-diamond distance with total Hamiltonian simulation time:
$\widetilde{\mathcal{O}}\left(C_{q}t_{\rm mix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\beta+1)}|\mathcal{A}|^{2}\log^{1+s}\left(1/\epsilon\right)\right)\,,$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{q}:=\log^{2+s}(S)+1\,.}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widetilde{\mathcal{O}}\left(\log(S^{2}+S\left\lVert H\right\rVert)+\log^{2}(t_{\rm mix}|\mathcal{A}|/\epsilon)+\log(\beta+2)\right)\,.}$
number of additional ancilla qubits for the prepare oracles and simulation. The $\widetilde{\mathcal{O}}$ absorbs a constant only depending on $s$, $\xi_{q}$, $\xi_{u}$, $\xi_{w}$, and $C_{1,u}$, and subdominant polylogarithmic dependencies on parameters $t_{\rm mix}$, $\|H\|$, $|\mathcal{A}|$, $S$, $\xi_{q}$, $\xi_{u}$, $\xi_{w}$, and $\beta$.
@@
3.4. Quasi-locality
It is worth noting that when $H$ is a local Hamiltonian and the coupling operators $\{A^{a}\}_{a\in\mathcal{A}}$ are local, then we can show that at sufficiently high temperature (small $\beta$), our jump operators $L_{a}$ in (3.3) are quasi-local, and the coherent term $G$ in (3.10) is a sum of quasi-local operators. This can be demonstrated by applying the Lieb-Robinson bounds [NS06, HK06, LR72, Has04, HHKL23]. Similar to [RFA24], for a finite lattice $\Lambda\subset\mathbb{Z}^{d}$, we consider a $(k,l)$-local Hamiltonian $H=\sum_{X\subset\Lambda}h_{X}$, which means that $h_{X}$ is a local potential with support on no more than $k$ sites, and each site appears in at most $l$ operators $h_{X}$.
Proposition 20 (Quasi-locality of jump and coherent operators).
Let $H=\sum_{X\subset\Lambda}h_{X}$ be a $(k,l)$-local Hamiltonian defined on a finite lattice $\Lambda$, and let each coupling operator $A^{a}$ be a local operator. The jump operator $L_{a}$ constructed via a Gaussian-type weighting function is quasi-local. Furthermore, there exists a constant $\beta_{c}>0$, which only depends on $k,l,\sup_{X}\|h_{X}\|$, such that for any $\beta<\beta_{c}$, the associated coherent operator $G$ is a sum of quasi-local operators.
Proof.
For simplicity, assume that each coupling $A^{a}$ is supported on a single site of the lattice denoted by $u_{a}$. We consider the Gaussian weighting function $q(\nu)=\exp(-(\beta\nu)^{2}/8)\omega(\nu/S)$ as the Gaussian-type one in (3.22) and set $S\geq 4\|H\|$. Define $\widehat{h}(\nu)=\exp(-(\beta\nu)^{2}/8)$. Recalling the formula for $L_{a}$ in (3.3), we have
(3.36) $L_{a}=\sum_{\nu\in B_{H}}\widehat{f}(\nu)A^{a}_{\nu}=\sum_{\nu\in B_{H}}\widehat{h}(\nu)A^{a}_{\nu}=\int^{\infty}_{-\infty}h(t)e^{iHt}A^{a}e^{-iHt}\,\mathrm{d}t\,,$
where we use $\omega(\nu/S)=1$ when $|\nu|\leq 2\|H\|$ in the second equality. Then, from the Lieb-Robinson bound [HHKL23, Lemma 5], there exists a constant $J$ that only depends on $k,l,\sup_{X}\|h_{X}\|$ such that
$\left\|e^{iHt}A^{a}e^{-iHt}-e^{iH_{B_{u_{a}}(r)}t}A^{a}e^{-iH_{B_{u_{a}}(r)}t}\right\|\leq\|A^{a}\|\frac{(J|t|)^{r}}{r!},$
for any $r>0$ and $a\in\mathcal{A}$. Here $B_{u_{a}}(r)\subset\Lambda$ is a ball centered at $u_{a}$ with radius $r$, and $H_{B_{u_{a}}(r)}=\sum_{X\cap B_{u_{a}}(r)\neq\emptyset}h_{X}$. This, along with (3.36), readily implies
(3.37) $\bigg{\|}L_{a}-\underbrace{\int_{-\infty}^{\infty}h(t)e^{iH_{B_{u_{a}}(r)}t}A^{a}e^{-iH_{B_{u_{a}}(r)}t}\,\mathrm{d}t}_{\text{\rm supported on }B_{u_{a}}(r)}\bigg{\|}$
$\leq$ $\frac{J^{r}\|A^{a}\|}{r!}\int_{-\infty}^{\infty}|h(t)||t|^{r}\,\mathrm{d}t=\|A^{a}\|\left(\mathcal{O}\left(J\beta/\sqrt{r}\right)\right)^{r}\,,$
for any $r>0$ and $\beta>0$, where the last equality is from $h(t)=\exp\left(-\Theta((t/\beta)^{2})\right)$. This implies that $L_{a}$ is quasi-local, and so is the operator $L_{a}^{\dagger}L_{a}$.
Define $\widehat{j}(\nu)=-\frac{i}{2}\tanh\left(-\beta\nu/4\right)$ with Fourier transform $j(t)$, defined in distribution sense, which decays as $\exp(-\Theta(|t|))$ as $|t|\to\infty$ and has singularity $\Theta(1/t)$ as $t\to 0$. We recall the definition (3.10) of $G$ and find
$\bigg{\|}G-\sum_{a\in\mathcal{A}}\underbrace{\int_{-\infty}^{\infty}j(t)e^{iH_{B_{u_{a}}(r)}t}(L_{a}^{\dagger}L_{a})e^{-iH_{B_{u_{a}}(r)}t}\,\mathrm{d}t}_{\text{\rm quasi local}}\bigg{\|}$
$\leq$ $\left(\sum_{a\in\mathcal{A}}\|L_{a}\|^{2}\right)\frac{J^{r}}{r!}\int_{-\infty}^{\infty}|j(t)||t|^{r}\mathrm{d}t=\left(\sum_{a\in\mathcal{A}}\|L_{a}\|^{2}\right)\left(\mathcal{O}\left(J\beta\right)\right)^{r}\,,\quad r\geq 1\,.$
This implies that setting $\beta_{c}=\Theta(J^{-1})$, $G$ is the sum of quasi-local operators if $\beta<\beta_{c}$. ∎
The above computation can be easily extended to the case where $q(\nu)=\widehat{h}(\nu)\omega(\nu/S)$ with $S\geq 4\|H\|$ and $|h(t)|\sim\exp(-\Theta(|t|^{p}))$ for $p\geq 1$ as $t\to\infty$. The quasi-locality does not affect the simulation complexity of the dynamics; however, it is a crucial property when analyzing the mixing time of the dynamics [RFA24]. We briefly address this in the following remark.
Remark 21.
To ensure that the constructed Lindblad dynamics (3.12) eventually relaxes to the desired Gibbs state, by 3, we should carefully choose the coupling operators $A^{a}$ and weighting functions $q^{a}$ such that the resulting $\{L_{a}\}$ and $G$ satisfy Eq. 2.5. This is always possible, due to the finite dimensionality of the system. Moreover, it is known [TKR+10, Theorem 9 and Section V] that for a primitive KMS detailed balanced QMS, the mixing time (1.3) can be estimated via the spectral gap of the Lindbladian. The recent work [RFA24] estimated the spectral gap of the efficient quantum Gibbs sampler in [CKG23] in the high-temperature regime, by mapping the Lindbladian $\mathcal{L}^{\dagger}_{\beta}$ with KMS DBC to a Hamiltonian ~$\widetilde{\mathcal{L}}_{\beta}^{\dagger}:=\sigma_{\beta}^{-1/4}\mathcal{L}^{\dagger}_{\beta}(\sigma_{\beta}^{1/4}X\sigma_{\beta}^{1/4})\sigma_{\beta}^{-1/4}$ and then analyzing its spectral properties by the stability of gapped Hamiltonians [MZ13]. A key fact for using the stability result [MZ13] is that the jump operators and the coherent term constructed in [CKG23] are quasi-local in the high-temperature regime. We prove in 20 that our $L_{a}$ and $G$ operators are also quasi-local at high temperatures, paving the way for future spectral gap analyses of the proposed Gibbs sampler.